OpenStudy (mrnebula):

HELP!!!!!!!!!!!!!!!!!! MEDAL AND FAN A rectangular prism has a width of 92 ft and a volume of 240 ft^2. Find the volume of a similar prism with a width of 46 ft. Round to the nearest tenth, if necessary. *30 ft^3 *40 ft^3 *60 ft^3 *120 ft^3

OpenStudy (mrnebula):

@severus4ever @undeadknight26

OpenStudy (campbell_st):

well when you compare the linear measurements the ratio is 46/92 of 1/2 then the volume will be \[V = (1/2)^3 \times 240 \]

OpenStudy (campbell_st):

the reason the scaling factor is cubed is because it's 3 dimensional if it was area it would be squared.... 2 dimensions... hope it helps

OpenStudy (mrnebula):

thnx